consider a car moving in the positive direction with an initial velocity of 39m/s slows down at a constant rate of -0.5m/s square. the distance traveled by car is ____ m
Answer #1

Using 3rd equation of motion,

V^2 = u^2 + 2as

0 = 39^2 - 2 x (-0.5) x s

Distance, s = 1521 m

Comment in case any doubt please rate my answer.....